# Zero-Shot Denoising for Fluorescence Lifetime Imaging Microscopy with Intensity-Guided Learning

### v1

@@ 4.3 Ablation Studies
As presented in Tab. 2 and illustrated in Fig. 4, each loss component contributes differently to FLIM denoising performance. The intensity loss ($\mathcal{L}_{intensity}$) alone effectively removes noise and achieves reasonable structural quality. However, due to the global optimization approach, it struggles to accurately preserve fluorescence lifetime information, resulting in a high Average Lifetime Error (ALE). Incorporating the fidelity loss ($\mathcal{L}_{fidelity}$) significantly enhances structural accuracy and lifetime consistency by enforcing signal-level coherence between input and reconstruction, although it insufficiently addresses noise reduction in the lifetime channels, as demonstrated in Fig. 4(c).
@@ 4.4 Discussion
While our framework currently demonstrates excellence in denoising fluorescence lifetime microscopy thanks to the real mixed Poisson-Gaussian noise datasets, its underlying methodology extends beyond this specific application. The proposed architecture can be seamlessly adapted to other multimodal microscopy techniques where similar noise patterns degrade visual and quantitative analysis.
Notably, our approach offers promising capabilities for super-resolution microscopy and deconvolution tasks, providing a computationally efficient solution that maintains exceptional performance. This approach requires minimal training data—a significant advantage in microscopy domains where annotated datasets are often limited.
By addressing the fundamental challenges of noise, resolution, and blur through a unified methodological approach, our framework represents a significant advancement in microscopy image enhancement technology with potential impact across multiple imaging modalities. These capabilities position our work as particularly relevant to quantitative biomedical imaging, where accurate downstream analysis depends on high-quality input data.
@@ 5 Conclusions
In this paper, we introduced a novel Zero-Shot Denoising Framework for FLIM with an Intensity-Guided Learning approach that effectively addresses the fundamental challenge of denoising multi-channel data in FLIM without requiring training examples for additional channels. By connecting fluorescence intensity and lifetime channels, our network significantly reduces noise while preserving critical lifetime information. Our comprehensive loss function design ensures effective denoising while maintaining the integrity of information across all channels. Experimental results on real-world datasets demonstrate that the proposed approach enables more reliable extraction of physiological and molecular information from challenging multimodal data acquisitions and potentially expands the applicability of this powerful technique to other related imaging techniques.

### v2

@@ 4.3 Ablation Studies
As presented in Tab. 2, each loss component contributes differently to FLIM denoising performance across all evaluated samples. The visual impact of these components is illustrated in Fig. 4 using the BPAE cell sample as a representative example. The intensity loss ($\mathcal{L}_{intensity}$) alone effectively removes noise and achieves reasonable structural quality. However, due to the global optimization approach, it struggles to accurately preserve fluorescence lifetime information, resulting in a high ALE. Incorporating the fidelity loss ($\mathcal{L}_{fidelity}$) significantly enhances structural accuracy and lifetime consistency by enforcing signal-level coherence between input and reconstruction, although it insufficiently addresses noise reduction in the lifetime channels, as specifically demonstrated in the BPAE cell sample shown in Fig. 4(c).
@@ 4.4 Discussion
The efficacy of our framework is dependent on the intensity prior, with enhanced performance directly yielding improved results across all lifetime channels. While our framework currently demonstrates excellence in denoising fluorescence lifetime microscopy, its underlying methodology extends beyond this specific application. The proposed architecture can be seamlessly adapted to other multimodal microscopy techniques with similar correlated noise patterns.
Notably, our approach offers promising capabilities for super-resolution microscopy and deconvolution tasks by providing a computationally efficient solution with exceptional performance. This method requires no training data —a significant advantage in microscopy domains where annotated datasets are often limited.
By addressing the fundamental challenges of noise, resolution, and blur through a unified methodological approach, our framework represents a significant advancement in microscopy image enhancement technology with potential impact across multiple imaging modalities. These capabilities establish our approach as highly applicable to quantitative biomedical imaging, where precise downstream analyses depend critically on high-quality input data.
@@ 5 Conclusions
In this paper, we introduced a novel Zero-Shot Denoising Framework for FLIM with an Intensity-Guided Learning approach that effectively addresses the fundamental challenge of denoising multi-channel data in FLIM without requiring training data for additional channels. By connecting fluorescence intensity and lifetime channels, our network significantly reduces noise while preserving critical lifetime information. The comprehensive loss function design ensures effective denoising while maintaining the integrity of information across all channels. Experimental results on real-world datasets demonstrate that the proposed approach enables more reliable extraction of physiological and molecular information from challenging multimodal data acquisitions and potentially expands the applicability of this powerful technique to other related imaging techniques.